OpenStudy (anonymous):

what is the function rule? x= -2, -1,0,1,2 y= 9, 4, -1, -6, -11

OpenStudy (freckles):

have you tried determine if it is a line or not?

OpenStudy (anonymous):

I don't understand what you mean?

OpenStudy (anonymous):

Are you talking about the x, y axis?

OpenStudy (freckles):

what i mean by line is have you tried to determine if your function is linear or not?

OpenStudy (freckles):

is your slope the same through out

OpenStudy (freckles):

\[m=\frac{y_2-y_1}{x_2-x_1} \text{ is this the same no matter what } (x_1,y_1) \text{ and } (x_2,y_2) \\ \text{ you choose from your table }\]

OpenStudy (freckles):

if that is then it is a line

OpenStudy (freckles):

and it will have form f(x)=mx+b

OpenStudy (anonymous):

I think it is linear because the is only a difference of 5 through, and I don't think it geometric

OpenStudy (freckles):

notice the x values you are given are evenly spaced and your y values form a arithmetic sequence that is there is a common difference so it will have a common slope amongst the points given therefore it is a linear function

OpenStudy (freckles):

so find the slope then replace m in the function rule: f(x)=mx+b with what you find the slope to be then use one of the points to find b

OpenStudy (anonymous):

y=-5

OpenStudy (freckles):

do you mean m=-5 ?

OpenStudy (freckles):

anyways I think you meant m=-5 \[f(x)=mx+b \\ f(x)=-5x+b\] so now you just need to find b use one of the points in plug them into the function rule f(x)=-5x+b And remember y=f(x)

OpenStudy (anonymous):

Well 9-5=4, and 4-5=-1 and -1-5= -6 so the function for y must be - 5

OpenStudy (anonymous):

y=f-5

OpenStudy (freckles):

I think you are getting the y coordinates confused with finding the slope clearly the y's aren't always -5 so you can't say y=-5

OpenStudy (anonymous):

oh..okay, so what does y equal then? I thought I was on the right track..

OpenStudy (freckles):

do you want to try to solve for b above given any of the points?

OpenStudy (freckles):

so far we have y=-5x+b I was asking you to use one of the points given to find b

OpenStudy (anonymous):

b=9??

OpenStudy (freckles):

what point did you use to determine that

OpenStudy (anonymous):

..I just used the y table...

OpenStudy (freckles):

(-2,9) is a point on your table (-1,4) is a point on your table (0,-1) is a point on your table (1,-6) is a point on your table (2,-11) is a point on your table you can use any of these (x,y) points to find b in the function rule f(x)=-5x+b

OpenStudy (freckles):

(0,-1) is actually the y-intercept if you didn't feel like doing any extra work

OpenStudy (anonymous):

Okay, I got it now, my teacher never explained that to me

OpenStudy (freckles):

f(x)=-5x+b b represents the y-intercept so b is?

OpenStudy (anonymous):

I don't know so I am guessing it's -11

OpenStudy (freckles):

y=mx+b m is the slope b is the y-intercept is this part familiar to you?

OpenStudy (anonymous):

No, my teacher wasn't here today to teach us this, so I am not sure

OpenStudy (anonymous):

I am sorry

OpenStudy (freckles):

ok we can just use one of the points given then y=-5x+b using (-2,9) to solve for b replace x with -2 and replace y with 9 9=-5(-2)+b can you try to solve this for b?

OpenStudy (anonymous):

is it 1?

OpenStudy (freckles):

close

OpenStudy (anonymous):

2!

OpenStudy (freckles):

further away -5 times -2 is 10

OpenStudy (freckles):

9=10+b

OpenStudy (freckles):

can you solve this for b now?

OpenStudy (freckles):

or still too hard?

OpenStudy (anonymous):

-1?

OpenStudy (freckles):

yes!

OpenStudy (freckles):

10+(-1) is totally 9 so b=-1

OpenStudy (freckles):

your function rule is f(x)=-5x-1

OpenStudy (freckles):

and see that matches are y-intercept (0,-1)

OpenStudy (freckles):

that b part does

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

I understand, now we just do something similar for the x axis

OpenStudy (freckles):

what

OpenStudy (freckles):

your job was to find the function rule we did that f(x)=-5x-1 is there something else you are suppose to do besides find function rule?

OpenStudy (anonymous):

No thank you

OpenStudy (freckles):

np

OpenStudy (anonymous):

You're very helpful!

OpenStudy (anonymous):

thank you a million times